OpenStudy (anonymous):

imogene's car traveled 280 miles averaging a certain speed. If the car had gone 5 mph faster the trip would have taken one hour less. Find the average speed

OpenStudy (anonymous):

t = d/r\[\left\{t=\frac{280}{r},t-1=\frac{280}{r+5}\right\} \]The following expression comes from eliminating t in the above list.\[r^2+5 r-1400=0 \]r = 35 mph